# Out-of-Distribution Segmentation in Autonomous Driving: Problems and State of the Art

### v1

@@ 3.5 Other
One of the main limitations of the methods that utilize outlier supervision is that the fine-tuning step decreases the performance of the original inlier model. In our own experiments, maximized entropy decreased the inlier performance by 1.6 mIoU points, and RbA decreased by 2.5 mIoU. In many practical scenarios, this is highly undesirable as OoD objects typically occur far less often than inlier objects. Instead of adjusting the inlier parameters of the model during the outlier supervision process, the authors of UEM [39] proposed an external module explicitly designed for OoD segmentation; this allows to keep the original inlier performance intact while performing very well in detecting OoD objects. The authors also suggested using the Likelihood Ratios between the learned distribution of outliers from the proxy OoD dataset and that of the original inlier model to calculate the OoD score. The authors found that using the feature representation of self-supervised pre-trained models like DinoV2 [43] resulted in better modelling of OoD objects than fully supervised models like Swin [31] or contrastive image-language pre-trained models like CLIP [45]
@@
In PixOOD [59], the authors propose an OoD segmentation method that does not require any outlier exposure. Their approach models the intra-class in-distribution variability through an online data condensation algorithm and consists of three components: 1) extracting the patch feature representation from a DinoV2 model, 2) building a 2D-projection space for each in-distribution class, and 3) finding calibrated in-distribution/OoD decision strategy. The framework builds on the methodology introduced in [58], but extends the approach for pixel-level predictions.
@@ 4 Limitations
Threshold Selection: One limitation of existing benchmarks for OoD segmentation is their reliance on threshold-free metrics (e.g., $AP$), which, while useful for assessing performance independent of a specific threshold, are not directly applicable to downstream tasks. In practice, a threshold must be selected to make reliable decisions, and ideally, it should effectively separate in-distribution and out-of-distribution regions with high confidence. Furthermore, the SMIYC-OT benchmark assesses component-level performance for practical reasons outlined in Section 2. However, their evaluation derives an optimal threshold from pixel-level metrics using the optimal pixel-level F1 score to assess component-level performance. This approach is troublesome for deployment scenarios, as ground-truth labels for OoD objects are often not available in large numbers for calibrating the OoD score, making it difficult to set a threshold in real-world settings. Moreover, this methodology obscures a critical limitation of some approaches: a narrow range of effective OoD scores. Some methods are highly sensitive to threshold selection, which means small variations can significantly impact their performance. This sensitivity is not adequately captured with current metrics, which may misrepresent the robustness of different approaches. Metrics like the Maximal Detection Margin [1] are a step in the right direction to mitigate some of these issues.

### v2

@@ 3.5 Other
One of the main limitations of the methods that utilize outlier supervision is that the fine-tuning step decreases the performance of the original inlier model. In our own experiments, maximized entropy decreased the inlier performance by 1.6 % mIoU points, and RbA decreased by 2.5 % mIoU. In many practical scenarios, this is highly undesirable as OoD objects typically occur far less often than inlier objects. Instead of adjusting the inlier parameters of the model during the outlier supervision process, the authors of UEM [42] proposed an external module explicitly designed for OoD segmentation; this allows to keep the original inlier performance intact while performing very well in detecting OoD objects. The authors also suggested using the Likelihood Ratios between the learned distribution of outliers from the proxy OoD dataset and that of the original inlier model to calculate the OoD score. The authors found that using the feature representation of self-supervised pre-trained models like DinoV2 [45] resulted in better modelling of OoD objects than fully supervised models like Swin [34] or contrastive image-language pre-trained models like CLIP [47]
@@
In PixOOD [61], the authors propose an OoD segmentation method that does not require any outlier exposure. Their approach models the intra-class in-distribution variability through an online data condensation algorithm and consists of three components: 1) extracting the patch feature representation from a DinoV2 model, 2) building a 2D-projection space for each in-distribution class, and 3) finding calibrated in-distribution/OoD decision strategy. The framework builds on the methodology introduced in [60] but extends the approach for pixel-level predictions. A similar direction is explored in [14], where the authors apply a k-nearest-neighbour-based strategy based on patch feature representations for dense OoD detection.
@@ 4 Limitations
Threshold Selection: One limitation of existing benchmarks for OoD segmentation is their reliance on threshold-free metrics (e.g., AUPRC), which, while useful for assessing performance independent of a specific threshold, are not directly applicable to downstream tasks. In practice, a threshold must be selected to make reliable decisions, and ideally, it should effectively separate in-distribution and out-of-distribution regions with high confidence. Furthermore, the SMIYC-OT benchmark assesses component-level performance for practical reasons outlined in Section 2. However, their evaluation derives an optimal threshold from pixel-level metrics using the optimal pixel-level F1 score to assess component-level performance. This approach is troublesome for deployment scenarios, as ground-truth labels for OoD objects are often not available in large numbers for calibrating the OoD score, making it difficult to select an optimal threshold in real-world settings. Moreover, this methodology obscures a critical limitation of some approaches: a narrow range of effective OoD scores. These methods are highly sensitive to threshold selection, which means small variations can significantly impact their performance. This sensitivity is not adequately captured with current metrics, which may misrepresent the robustness of different approaches. Metrics like the Maximal Detection Margin [1], are a step in the right direction to mitigate some of these issues.
@@ 5 Future Directions
In addition to detection, context-aware decision-making is essential: not all OoD objects pose the same level of risk. Thus, it becomes crucial to interpret the OoD detection in light of the surrounding scene context and possible future interactions. This requires integrating scene understanding and behaviour prediction with the OoD segmentation pipeline. For instance, distinguishing between a static roadside billboard and a moving pedestrian in an unknown outfit directly influences the vehicle’s response. Such scene-level understanding supports more adaptive and risk-sensitive planning strategies in real-world applications.